# RetrySQL: text-to-SQL training with retry data for self-correcting query generation

### v1

@@ 4.2. Linear probing dataset
4.3. Supervised fine-tuning with LoRA
In addition to continued pre-training, we also examined the scenario of SFT with retry data, reflecting a practical situation where an open-source pre-trained model is adapted to enhance reasoning capabilities or to address a specific task. Our objective was to evaluate whether fine-tuning with retry data is as effective as continued pre-training. We focused on PEFT, specifically LoRA (Hu et al., 2022), which is a technique commonly used in practice. LoRA fine-tunes a limited set of trainable parameters, keeping the original pre-trained weights frozen. Typically, for Transformer-based dense models, LoRA is applied either to just the query/value matrices or to all linear layers of the network (Hu et al., 2022; Mao et al., 2024). Since in the fine-tuning process we added new special tokens (namely [REASONING] and [BACK]), both input and output embeddings needed to be trained as well. We fine-tuned the base model with retry data directly, we did not start with a model pre-trained using only error-free data (contrary to (Ye et al., 2024b)). In the latter configuration, LoRA is incapable of leveraging the retry data without destroying the previously learned knowledge.
4.4. Inference process and evaluation metrics
For the purpose of measuring the effectiveness of retry data in text-to-SQL generation, we utilized the Execution Accuracy (EX) metric introduced in the BIRD benchmark. In all experiments we used the development dataset provided by BIRD for evaluation. It contains a total of 1534 examples. The data format is the same as the training set, with the addition of a difficulty value (simple, moderate, challenging) for each example. During inference, we used the same format as during training, but omitted the part of each sequence after the [REASONING] token. Thus, we wanted the model to first generate the reasoning steps, and then the [SQL] token, followed by the actual SQL query.
During inference we used the best checkpoint of each trained model variant (except for SFT with LoRA, where we used the last checkpoint, following  (Ye et al., 2024b)). We used beam search multinomial sampling with 4 beams as a decoding strategy (following (Ye et al., 2024b)), with $temperature=0.5$, $top\_k=50$ and $top\_p=1.0$. We limited the number of new tokens to 1024. Each evaluation example was processed 5 times, for the purpose of measuring the model variance. The results were post-processed by removing the [SQL] and preceding tokens, leaving only the SQL query.
We then calculated the EX metric in the following manner. First, the generated SQL query as well as the ground truth query for a given question were executed in an SQLite database containing the development data. Then, resulting sets of rows were compared and the ratio of matching rows was saved. Finally, the match ratios for all examples were averaged. We report four EX values: EXoverall, EXsimple, EXmoderate, EXchallenging. They correspond to an overall EX over all examples, or over just the simple, moderate or challenging ones, respectively.
5. Results
In this section, we assess the effectiveness of RetrySQL for training SQL generation models. We showcase the baseline results for both proprietary and open-source LLMs (Section 5.1). We demonstrate through linear probing that the baseline OpenCoder model can recognize incorrect reasoning steps (Section 5.2). We then show that training with retry data improves Execution Accuracy compared to training with error-free reasoning steps (Section 5.3). We verify that SFT with LoRA is incapable of training models that utilize retry data (Section 5.4). Finally, we explain the self-correction behavior with an analysis of model confidence around the [BACK] tokens (Section 5.5).
5.1. Zero-shot baselines
We evaluated OpenCoder 1.5B , as well as several proprietary LLMs, in zero-shot mode (Tab. 1). We observe that OpenCoder 1.5B falls behind state-of-the-art models, with an EXoverall score gap ranging from 7.95 percentage points (p.p.) for GPT-4o-mini, to 39.31 p.p. for Gemini-1.5-pro. The same is true for the detailed metrics as well, which overall signifies relatively poor performance in the text-to-SQL task. These results indicate that, in the zero-shot setting, even a recent open-source coding-oriented model such as OpenCoder is not able to reach the performance of general-purpose proprietary solutions.
5.2. Detecting regretful patterns through linear probing
Figure 2. t-SNE projection of OpenCoder’s internal state embeddings for the linear probing task. Blue points represent embeddings corresponding to correct reasoning steps, while orange points indicate embeddings for incorrect steps. The clusters of orange points indicate that the OpenCoder model differentiates a large portion of the incorrect steps from the correct ones, highlighting the innate, yet hidden, ability to detect mistakes in the reasoning process.
We took the baseline OpenCoder model and performed a linear probing experiment with its frozen weights. The linear probing model detected incorrect steps with average $balanced\_accuracy=82\%$ and $f1\_score=71\%$. Since the detection accuracy was significantly higher than $50\%$ (random guess), we can conclude that the probing signals most likely came from the pre-trained weights, and not from the fine-tuned classification layer (Ye et al., 2024a). To support our findings, we visualize the internal state embeddings of OpenCoder using a t-SNE projection (Fig. 2). The plot illustrates the separability of the model’s internal states when it comes to predicting correct versus incorrect reasoning steps. These results demonstrate that OpenCoder has an innate capability to self-correct, which is in line with previous research positing that this ability is a universal law for all Transformer models (Ye et al., 2024b). This justifies the usage of retry data, which should enable the model to backtrack as it generates an incorrect step and then retry once again.
Table 2. Execution Accuracy for the continued pre-training case with RetrySQL. All results are expressed in percentages, with mean and standard deviation over 5 multinomial beam search generations. The best results are marked in bold. Results with retry data that improve upon the error-free training are indicated with an underline.
Dataset variant EXsimple EXmoderate EXchallenging EXoverall
– (zero-shot) 40.04 ± 0.20 16.90 ± 0.48 7.45 ± 0.91 29.96 ± 0.07
error-free 62.70 ± 0.07 43.53 ± 0.14 39.45 ± 0.28 54.71 ± 0.08
Retry FS 0.1 65.84 ± 0.00 44.09 ± 0.11 36.83 ± 0.34 56.52 ± 0.04
Retry FS 0.2 68.22 ± 0.12 45.47 ± 0.14 40.28 ± 0.34 58.70 ± 0.09
Retry FS 0.3 68.00 ± 0.00 44.91 ± 0.26 43.31 ± 0.28 58.68 ± 0.06
Retry FS 0.4 66.57 ± 0.11 44.22 ± 0.32 34.62 ± 0.28 56.79 ± 0.11
Retry FS 0.5 66.98 ± 0.18 44.96 ± 0.29 37.79 ± 0.28 57.56 ± 0.12
5.3. Retry data improves SQL generation metrics
To test if the OpenCoder model can learn the ability to self-correct, we continued the pre-training process with reasoning-enhanced BIRD training data, both error-free and with retry data. Compared to the zero-shot OpenCoder 1.5B baseline, the model continuously pre-trained with error-free data yielded an impressive improvement in generation accuracy metrics (Tab. 2): $\sim$23 p.p. for simple examples, $\sim$27 p.p. for moderate ones, and 32 p.p. for challenging instances, which resulted in an overall increase of $\sim$25 p.p.. These results are expected, as during training we provided the model with previously unseen domain-specific text-to-SQL training samples.
The retry data results show us that models continuously pre-trained with such corrupted samples lead to improved accuracy metrics when compared to the error-free continued pre-training (Tab. 2, Tab. S1). Out of four approaches to retry data preparation, the FS variant proved to be the best overall. Thus, here we present findings only for the FS variant. For the results pertaining to the other variants (i.e. FM, FBS, FBM), see Appendix A.2. The highest improvement in overall generation accuracy, 3.99 p.p., was observed for $p_{retry}$=0.2. The second best increase, 3.97 p.p., resulted from employing the Retry FS 0.3 dataset.
While the EXoverall metric does show the effectiveness of using retry data for improving text-to-SQL generation accuracy in general, it does not show the full picture. Looking at the EXsimple score, we see that the biggest improvement stems from using the dataset with $p_{retry}$=0.2 (increase of 5.52 p.p.). For the moderate cases, the biggest improvement of the EX score equals 2.15 p.p. (once again for $p_{retry}$=0.2). The biggest increase in the EXchallenging metric, 4.00 p.p., resulted from pre-training with the Retry FS 0.3 dataset. Previous research on self-correcting generation with retry data postulated that the advantage of the self-correction ability can be observed especially for complex out-of-distribution evaluation examples, which required the longest solution reasoning sequences (Ye et al., 2024b). Our results show that for the text-to-SQL task, the improvements are more even across difficulty levels. This is a consequence of how the BIRD dataset is constructed and of our approach to generating reasoning steps (see Section 3.2) - the number of operations required to explain SQL queries is not perfectly correlated with the difficulty level. There is a significant overlap in the complexity of SQL queries across difficulty levels in the BIRD development dataset (Fig. S1). As such, many challenging examples can be solved with relatively short queries, which do not necessitate the model to generate long reasoning step sequences. Examples that do require more reasoning are simply spread out across all difficulty levels, which explains the relatively even increases of Execution Accuracy observed in our experiments. These results show that the self-correction ability introduced by pre-training with retry data is an effective way of improving the accuracy of SQL generation.
Table 3. Execution Accuracy for SFT with LoRA. We performed LoRA fine-tuning using only the Retry FS 0.3 variant. All results are expressed in percentages, with mean and standard deviation over 5 multinomial beam search generations.
LoRA rank EXsimple EXmoderate EXchallenging EXoverall
– (error-free) 62.70 ± 0.07 43.53 ± 0.14 39.45 ± 0.28 54.71 ± 0.08
${r=8}$ 14.98 ± 0.20 7.07 ± 0.32 5.24 ± 0.34 11.67 ± 0.17
${r=16}$ 56.22 ± 0.27 30.34 ± 0.21 23.72 ± 0.70 45.32 ± 0.17
${r=32}$ 56.76 ± 0.22 30.60 ± 0.49 22.62 ± 0.28 45.62 ± 0.15
${r=64}$ 60.02 ± 0.16 36.38 ± 0.16 30.34 ± 0.00 50.07 ± 0.09
${r=128}$ 58.90 ± 0.16 36.98 ± 0.17 23.03 ± 0.34 48.88 ± 0.09
${r=256}$ 58.96 ± 0.09 38.45 ± 0.17 28.55 ± 0.34 49.88 ± 0.12
5.4. SFT with LoRA is ineffective for training with retry data
For the experiments that evaluated SFT with LoRA, we used the Retry FS 0.3 dataset as the training data. While it scored marginally lower than the Retry FS 0.2 variant in the overall EX metric, the high performance for the challenging examples marked it as a preferable dataset.
It is evident that SFT with LoRA is not effective when it comes to learning the self-correction ability from retry data (Tab. 3). None of the tested LoRA ranks resulted in models that improved upon the error-free training. For rank 8 the model was not able to learn almost anything from the training data ($EX_{overall}$ equal to 11.67%). For higher ranks some knowledge was transferred (overall accuracy up to 50.07%), but the limited number of trainable parameters was too low to effectively preserve both the reasoning step generation and self-correction abilities at the same time.
Consequently, to utilize the retry data and teach a coding LLM to self-correct in the text-to-SQL task, it is necessary to perform full-parameter continued pre-training. This finding supports observations made in previous work regarding the capability of LoRA-trained models to learn from retry data (Ye et al., 2024b).
While the usage of LoRA is the only way for most practitioners to fine-tune sizable (70B parameters and more) LLMs, the fact that a small 1.5B-parameter model trained with RetrySQL achieves very good results indicates that continued pre-training of smaller models with the addition of reasoning-based retry data might be a viable alternative.
Figure 3. Distribution of token confidence before and after [BACK] tokens. (a) Mean of max token confidence across 10 beam search passes. It can be seen that the confidence score is on average much higher for tokens after the [BACK] token, indicating that the model is uncertain as it makes mistakes, but is confident after self-correction. (b) Standard deviation of max token confidence across 10 beam search passes. The variance of model predictions is much higher as it makes mistakes than after self-correction. Both of these results show that the self-correcting behavior is indeed learned by a model trained with RetrySQL. We used RetrySQL-FS-0.3 for obtaining these results.
5.5. Does RetrySQL know that it makes mistakes?
The EX results show us that training with RetrySQL increases the number of correct SQL queries compared to models trained without retry data (Tab. 2). However, it could be the case that the additional tokens present in reasoning steps in retry data simply act as robust augmentation and the model does not learn to self-correct. The underlying model behavior needs to be studied in more detail. To this end, we analyzed the confidence scores returned by the model trained with the Retry FS 0.3 dataset (denoted further as RetrySQL-FS-0.3) in proximity (radius of 10 tokens) to [BACK] tokens. We took max softmax scores from these tokens, and then calculated the mean and standard deviation per token across 10 multinomial beam search passes. Finally, we averaged these metrics separately for tokens before and after each [BACK] token.
Both of these results clearly show that the self-correcting ability is a learned behavior, resulting from the inclusion of retry data in the training process. For examples of model output that contains reasoning steps with self-correction, see Appendix A.7.
Table 4. Execution Accuracy for the full end-to-end text-to-SQL pipeline. All results are expressed in percentages, with mean and standard deviation over 5 multinomial beam search generations.
Model name EXsimple EXmoderate EXchallenging EXoverall
GPT-4o 61.62 ± 3.83 42.74 ± 2.34 40.48 ± 1.78 54.99 ± 0.32
GPT-4o-mini 42.12 ± 1.06 18.06 ± 0.47 17.65 ± 0.38 32.53 ± 0.72
Gemini-1.5-pro 66.88 ± 0.23 48.15 ± 0.62 51.86 ± 0.58 59.79 ± 0.31
Gemini-1.5-flash 64.69 ± 0.12 45.86 ± 0.24 49.52 ± 0.57 57.56 ± 0.08
RetrySQL-FS-0.2 59.81 ± 0.04 37.46 ± 0.09 35.17 ± 0.00 50.72 ± 0.00
RetrySQL-FS-0.3 60.28 ± 0.05 38.36 ± 0.14 36.00 ± 0.28 51.36 ± 0.05
5.6. RetrySQL-trained model in an end-to-end text-to-SQL pipeline
All experiments presented thus far focused strictly on the generation step, with perfect pre-computed schema linking. In that setting, we found that the best RetrySQL-trained models, denoted further as RetrySQL-FS-0.2 and RetrySQL-FS-0.3 (trained with Retry FS 0.2 and Retry FS 0.3, respectively), achieved EXoverall scores of 58.70% and 58.68% (Tab. 2). These metrics are significantly higher than the result for the proprietary GPT-4o-mini model (by $\sim$20.8 p.p.), and quite close to the score for the GPT-4o-model (short of it by $\sim$5 p.p.) (Tab. 1). However, in a real text-to-SQL pipeline, perfect schema linking is not available. In order to validate if RetrySQL-FS models are competitive with existing proprietary models in the full end-to-end pipeline setting, we performed an additional set of experiments. For the complete description of the pipeline, together with details on our schema linking methodology, see Appendix A.3.
In the full pipeline setting, we observed that the RetrySQL-trained models remain competitive with much larger models (Tab. 4). The EXoverall scores for RetrySQL-FS-0.2 and RetrySQL-FS-0.3 are still much higher than the result for GPT-4o-mini (by $\sim$18-19 p.p.). At the same time, the gap to GPT-4o is smaller in the full pipeline setting: 3.63 p.p. (for RetrySQL-FS-0.3), compared to $\sim$5 p.p. for the perfect schema linking setting.
Crucially, it must be highlighted that the RetrySQL-FS models are relatively small (1.5B parameters). The aforementioned proprietary GPT-4o-mini and GPT-4o LLMs are estimated to be much larger ($\sim$8B and $\sim$200B parameters (Abacha et al., 2025), respectively).
These results indicate that using retry data in conjunction with our RetrySQL method produces 1.5B-parameter models that are competitive with much larger proprietary models, as measured by the BIRD EX metric. This is a promising outcome, showing that incorporating self-correction in the generation stage might be a way forward for future text-to-SQL end-to-end pipelines.
6. Limitations
In our experiments we used the training data from the BIRD benchmark, which contains a limited number of training examples. This is different than what has been done in previous work on self-correction with retry data (Ye et al., 2024b). There, the training examples were generated on demand as the training went on, to fill a preset number of training steps. We did not have a setup for generating synthetic data in that way, and had to rely on the curated training examples from BIRD. This might explain the discrepancy in the relative effectiveness of retry data - with the on-demand synthetic examples, the observed improvements in generation accuracy were in the range of 10-16 p.p. for the hardest examples. However, it is important to keep in mind that a direct comparison to our results is not obvious, as the problem setting of grade school math is very different to our text-to-SQL task. In addition, the metric used in that work was a direct measure of correctness, while in our case we used an indirect Execution Accuracy metric computed in relation to database values, which were not present in the training data. Furthermore, as indicated previously (see Section 5.3), the difficulty of examples in the BIRD development set does not correlate with reasoning length. However, despite these differences, the effectiveness of our approach is still evident, as the EX metric was noticeably improved thanks to the addition of retry data in the training examples. We leave synthetic training data generation as a topic for future work.
7. Conclusions
In this paper, we presented RetrySQL, a novel approach to training text-to-SQL generation models. Our solution utilizes reasoning steps with retry data in the training examples, which teaches the generation model to self-correct itself as it produces its output. We show that using such data for the continued pre-training of a coding LLM leads to improved Execution Accuracy metrics when compared to models pre-trained without retry data (increase of $\sim$4 p.p. both overall and for the challenging examples). In addition, we confirm previous observations related to using SFT with LoRA for the purpose of training with retry data. We provide an explainability analysis for our results - we show that as the RetrySQL model makes mistakes, it is less confident in its predictions than after it self-corrects itself. Finally, we showcase that incorporating RetrySQL-trained 1.5B-parameter models into a relatively simple end-to-end text-to-SQL pipeline produces results that are competitive with much larger closed-source proprietary LLMs such as GPT-4o-mini and GPT-4o. We hope that our RetrySQL training paradigm will lead to further developments in text-to-SQL models, especially in the context of self-correcting generation.
Appendix A Appendix
A.1. Training details
All models were trained with NVIDIA A100 80GB, utilizing 2 GPUs. The effective batch size equaled 128, due to gradient accumulation being set to 4 and the per-device batch size set to 16. We used dynamic per-batch right padding, with sequence length being padded to a multiple of 8. We utilized the AdamW optimizer with the following hyperparameters (inspired by (Li et al., 2024)): $\beta_{1}$=0.9, $\beta_{2}$=0.95, $epsilon$=1e-08, $weight\_decay$=0.1. In all experiments, we trained for 5 epochs, with $learning\_rate$=5.0e-05 and a cosine learning rate schedule. For optimizing the memory usage, we employed the DeepSpeed framework (Rasley et al., 2020) with stage Zero-2.
For experiments regarding SFT with LoRA, we utilized the same hardware and hyperparameters as described above. Following previous work (Ye et al., 2024b), we examined this set of low-rank configurations: ${r\in\{8,16,32,64,128,256\}}$. We applied $lora\_dropout=0.01$ and set $\alpha=2r$, following the common best practices (Hu et al., 2022).
A.2. Full RetrySQL results
In this section, we provide a full evaluation of all retry data variants (Tab. S1), together with a visualization for the distribution of SQL query complexity across data instances in the BIRD development dataset (Fig. S1).
@@ A.2. Full RetrySQL results
Table S1. Execution Accuracy for the continuous pre-training case with RetrySQL, with all retry data variants. The OpenCoder 1.5B model was used as the starting point for all trainings. All results are expressed in percentages. Since we found that the model variance across multinomial beam search passes is relatively low for the FS datasets (Tab. 2), we did not calculate standard deviations for the remaining variants. The best results are marked in bold. Results with retry data that improve upon the error-free training are indicated with an underline.
@@
error-free 62.70 43.53 39.45 54.71
@@
Figure S1. Distribution of SQL query complexity in the BIRD development dataset. There is a significant overlap in the number of level-1 expressions in the SQL syntax tree across difficulty levels defined in the BIRD development dataset. Due to our reasoning generation strategy (see Section 3.2), the number of these expressions is a proxy for the number of reasoning steps. We parsed the ground truth SQL queries with the SQLGlot Python library (SQLGlot, [n. d.]) and extracted level-1 elements from the corresponding syntax trees.
A.3. Text-to-SQL pipeline
@@ A.3.1. Pipeline description
A.3.2. Schema linking
Schema linking is one of the most critical steps within the full text-to-SQL pipeline. While recent advancements suggest that it might be omitted in the case of using LLMs for the query generation (Maamari et al., 2024), feeding the model with too much data can provoke hallucinations and increase the inference cost due to a large number of processed input tokens (Chung et al., 2025). Moreover, the full database schema might simply not fit into the context window of the model.
Inspired by leading solutions in the BIRD benchmark (Pourreza and Rafiei, 2023, 2024), we treated schema linking as a separate task and designed a bespoke experimentation framework solely for the purpose of evaluating selected schema linking solutions. We extracted ground truth schema links from the BIRD development set and then investigated several schema linking algorithms. The tested algorithms can be grouped in the following categories:
@@ A.3.2. Schema linking
Table S2. Results for the schema linking experiments. The best performing approach, as measured by all metrics, was the LLM-based schema linking with Gemini-1.5-pro. For FP, lower values are better. For recallcol and recalllink higher values are better. The best results are marked in bold.
@@
Snowflake - arctic-embed-m (Merrick, 2024) 0.75 0.6 0.23
Snowflake - arctic-embed-l (Merrick, 2024) 0.89 0.78 0.48
@@
A.4. Linear probing experiments
In this section we provide more details regarding our linear probing experiments, which were inspired by a probing task originally introduced in previous work (Ye et al., 2024a) (specifically, the $can\_next(A)$ task).
A.4.1. Experiment setup
We prepared the linear probing model by taking the OpenCoder-1.5B weights continuously pre-trained with error-free reasoning steps and replacing the existing head with a linear one. The new head mapped the $2044$-dimensional vector of the last token in the input sequence to a single sigmoid-activated neuron for the binary classification task. Unlike (Ye et al., 2024a), we did not introduce any small rank-r update on the input (embedding) layer. We trained this classification model using a machine equipped with 2 x NVIDIA A100 80GB, with effective $batch\_size$=128 and $learning\_rate$=1e-4. We used the AdamW optimizer with $\beta_{1}$=0.9, $\beta_{2}$=0.95 and utilized early stopping.
We trained the model in a 5-fold cross-validation setup, by retaining $80\%$ of the dataset for training and using the rest for validation. We used the best checkpoints to compute $balanced\_accuracy$ and $f1\_score$ (reported for the incorrect class).
A.4.2. Discussion
By performing the linear probing experiment we demonstrated that a coding-oriented LLM, previously pre-trained with only error-free reasoning steps, exhibits regretful patterns during inference for input examples that contain an incorrect sequence of reasoning steps. It is possible to leverage the probing results by applying the classifier described above to guide the SQL generation process. After generating each solution sentence, the next probing could be used to determine whether the model knows that it has made a mistake and, if so, it could be reverted to the end of the previous sentence and regenerate from that point. Previous work showed that this method can increase generation accuracy in mathematical reasoning tasks, at the cost of increased inference complexity (Ye et al., 2024b). We did not perform such an analysis, since we treated the probing experiment as a preliminary sanity check for using retry data in the text-to-SQL task.
A.5. Training data example

### v2

@@ A.2 Zero-shot evaluation results
A.3 Full RetrySQL results
In this section, we provide a full evaluation of all retry data variants (Tab. S2) for OpenCoder 1.5B and BIRD, as well as several additional analyses for the results.
First, we assessed the impact of adding error-free reasoning steps to BIRD training data. It is evident that they have a significant impact on model performance by themselves - the overall Execution Accuracy increases by over 17 p.p. compared to training without reasoning steps. However, with RetrySQL it is possible to increase the model performance even further by enabling the resulting model to self-correct itself during generation.
@@ A.3 Full RetrySQL results
Table S2: Execution Accuracy for OpenCoder 1.5B continuously pre-trained with RetrySQL and evaluated on BIRD, with all retry data variants. All results are expressed in percentages. Since we found that the model variance across multinomial beam search passes is relatively low for the FS datasets (Tab. 1), we did not calculate standard deviations for the remaining variants. The best results are marked in bold. Results with retry data that improve upon the error-free training are indicated with an underline.
@@
error-free (no reasoning) 43.78 28.88 24.83 37.48
error-free (with reasoning) 62.70 43.53 39.45 54.71
@@
Figure S1: Distribution of SQL query complexity in the BIRD development dataset. There is a significant overlap in the number of level-1 expressions in the SQL syntax tree across difficulty levels defined in the BIRD development dataset. Due to our reasoning generation strategy (see Section 3.2), the number of these expressions is a proxy for the number of reasoning steps. We parsed the ground truth SQL queries with the SQLGlot Python library (sqlglot) and extracted level-1 elements from the corresponding syntax trees.
A.4 Additional result analysis
We performed several additional analyses to explain the behavior of RetrySQL. You can find their outcomes below:
  • Query complexity We plotted the distribution of SQL query complexity across data instances in the BIRD development dataset (Fig. S1). There is a significant overlap of query complexity for all difficulty levels.
  • Number of [BACK] tokens per difficulty We measured the mean number of generated [BACK] tokens per difficulty level for responses generated by a RetrySQL-trained model (OpenCoder 1.5B, Retry FS 0.2, BIRD dataset): simple 1.15 ± 0.35, moderate 1.46 ± 1.02, challenging 1.57 ± 0.73. There is no clear correlation in the mean number of [BACK] tokens per difficulty level.
  • Reasoning chain length for error-free and RetrySQL-trained models The average number of steps for the error-free model is 12.48 ± 7.73, while for a model trained with Retry FS 0.3 it is 9.472 ± 4.18 (Tab. S3). The reason for the lower number of steps with the RetrySQL-trained model comes from the presence of [BACK] tokens, i.e. from the self-correcting ability. The error-free model doesn’t have any means to self-correct, but it knows when a mistake was made (see Section 5.2). The only way for it to attempt a correction is to generate more reasoning steps (Fig. S11). On the other hand, the RetrySQL-trained model does have the self-correcting ability, so whenever it detects a mistake, it simply generates a [BACK] token and corrects it, which leads to an overall reduction in the number of required reasoning steps.
Table S3: Mean number of reasoning steps generated by OpenCoder 1.5B. We measured the number of reasoning steps for examples from the BIRD evaluation dataset. Error-free model was trained without retry data, while RetrySQL was trained with Retry FS 0.3.
Model variant Simple Moderate Challenging Overall
All evaluation examples
error-free 11.71 ± 7.67 11.82 ± 7.29 13.62 ± 8.88 12.48 ± 7.73
RetrySQL 8.638 ± 4.17 9.91 ± 4.40 9.754 ± 2.97 9.472 ± 4.18
Evaluation examples with generated [BACK] token
error-free 14.516 ± 9.77 13.26 ± 8.62 16.29 ± 10.52 14.34 ± 9.57
RetrySQL 10.81 ± 5.19 12.556 ± 6.03 10.57 ± 1.59 11.39 ± 5.20
Evaluation examples without generated [BACK] token
error-free 10.93 ± 6.79 11.06 ± 6.38 12.23 ± 6.79 11.06 ± 6.69
RetrySQL 8.04 ± 3.62 8.55 ± 2.34 9.59 ± 3.78 8.30 ± 3.37
A.5 Text-to-SQL pipeline
@@
Schema linking
Schema linking is one of the most critical steps within the full text-to-SQL pipeline. While recent advancements suggest that it might be omitted in the case of using LLMs for the query generation (maamari_death_2024), feeding the model with too much data can provoke hallucinations and increase the inference cost due to a large number of processed input tokens (chung_longcontext_2025). Moreover, the full database schema might simply not fit into the context window of the model.
Inspired by leading solutions in the BIRD benchmark (pourreza_din-sql_2023; pourreza_dts-sql_2024), we treated schema linking as a separate task and designed a bespoke experimentation framework solely for the purpose of evaluating selected schema linking solutions. We extracted ground truth schema links from the BIRD development set and then investigated several schema linking algorithms. The tested algorithms can be grouped in the following categories:
@@
Table S4: Results for the schema linking experiments. The best performing approach, as measured by all metrics, was the LLM-based schema linking with Gemini-1.5-pro. For FP, lower values are better. For recallcol and recalllink higher values are better. The best results are marked in bold.
@@
Snowflake - arctic-embed-m (arctic_embed) 0.75 0.6 0.23
Snowflake - arctic-embed-l (arctic_embed) 0.89 0.78 0.48
@@
Pipeline results
Here we provide Execution Accuracy results for the full pipeline setting, evaluated with the BIRD benchmark dataset. RetrySQL-trained OpenCoder 1.5 models compare favorably to much larger, proprietary models: they achieve EXoverall of up to 51.36, compared to 32.53 and 54.99 for GPT-4o-mini and GPT-4o. Crucially, they still improve over error-free training in this case as well, which shows that RetrySQL remains effective even with non-perfect schema linking.
Table S5: Execution Accuracy for the full end-to-end text-to-SQL pipeline. All results are expressed in percentages, with mean and standard deviation over 5 multinomial beam search generations.
Model name EXsimple EXmoderate EXchallenging EXoverall
GPT-4o 61.62 ± 3.83 42.74 ± 2.34 40.48 ± 1.78 54.99 ± 0.32
GPT-4o-mini 42.12 ± 1.06 18.06 ± 0.47 17.65 ± 0.38 32.53 ± 0.72
Gemini-1.5-pro 66.88 ± 0.23 48.15 ± 0.62 51.86 ± 0.58 59.79 ± 0.31
Gemini-1.5-flash 64.69 ± 0.12 45.86 ± 0.24 49.52 ± 0.57 57.56 ± 0.08
error-free 55.95 ± 0.09 38.94 ± 0.09 31.03 ± 0.0 48.47 ± 0.07
RetrySQL-FS-0.2 59.81 ± 0.04 37.46 ± 0.09 35.17 ± 0.00 50.72 ± 0.00
RetrySQL-FS-0.3 60.28 ± 0.05 38.36 ± 0.14 36.00 ± 0.28 51.36 ± 0.05
A.6 Linear probing experiments
In this section we provide more details regarding our linear probing experiments, which were inspired by a probing task originally introduced in previous work (ye_physics_part2point1_2024) (specifically, the can_next(A)can\_next(A) task).
Experiment setup
We prepared the linear probing model by taking the OpenCoder-1.5B weights continuously pre-trained with error-free reasoning steps and replacing the existing head with a linear one. The new head mapped the $2044$-dimensional vector of the last token in the input sequence to a single sigmoid-activated neuron for the binary classification task. Unlike (ye_physics_part2point1_2024), we did not introduce any small rank-r update on the input (embedding) layer. We trained this classification model using a machine equipped with 2 x NVIDIA A100 80GB, with effective batch_sizebatch\_size=128 and learning_ratelearning\_rate=1e-4. We used the AdamW optimizer with $\beta_{1}$=0.9, $\beta_{2}$=0.95 and utilized early stopping.
We trained the model in a 5-fold cross-validation setup, by retaining $80\%$ of the dataset for training and using the rest for validation. We used the best checkpoints to compute balanced_accuracybalanced\_accuracy and f1_scoref1\_score (reported for the incorrect class).
Discussion
By performing the linear probing experiment we demonstrated that a coding-oriented LLM, previously pre-trained with only error-free reasoning steps, exhibits regretful patterns during inference for input examples that contain an incorrect sequence of reasoning steps. It is possible to leverage the probing results by applying the classifier described above to guide the SQL generation process. After generating each solution sentence, the next probing could be used to determine whether the model knows that it has made a mistake and, if so, it could be reverted to the end of the previous sentence and regenerate from that point. Previous work showed that this method can increase generation accuracy in mathematical reasoning tasks, at the cost of increased inference complexity (ye_physics_part2point2_2024). We did not perform such an analysis, since we treated the probing experiment as a preliminary sanity check for using retry data in the text-to-SQL task.
A.7 Training data example